1) 20 women went to the market to buy a bag of rice which cost 50cedis per bag. 10 women had 60cedis each and 5 women had 40cedis each and the rest of the women had 25cedis each. Calculate :


i) The level of demand

ii) Want or desire want


1
Expert's answer
2021-03-02T07:57:29-0500

Solution:

i). The level of demand:

a bag of rice = 50 cedis

Total amount the 20 women have ="(10\times 60) +(5\times 40)+(5\times 25)"

= 600 + 200 +125 = 925 cedis

The level of demand = "\frac{Total\; amount}{price \;of \;a \;bag\; of\; rice}"


="\frac{925}{50}" = 18.5 bags

The level of demand = 19 bags of rice

 

ii). Want or desire want:

A want is a desire to have something. All 20 women want rice but only 10 women can afford it.

The want = The difference required to buy the rice

 = "(5\times 10) +(5\times 25)" = 50 + 125 = 175 cedis


Want or desire want = 175 cedis